Given f (x)= x2-10x/x3-x+20, find f (-1)
Answer by checkley77(12844)

f (x)= x^2-10x/x^3-x+20,
find f (-1)
f(-1)=-1^2-10*-1/-1^3-(-1)+20
f(-1)=1+10/-1+1+20
f(-1)=11/20 ans.